A bullet of mass m moving with a speed v strikes a wooden block of mass M and gets embedded into the block. Find the final speed. Answer is root (M/M+m)v

Grade:9

1 Answers

Himani sharma
5 years ago
It can be solved by consevation of momentum i e m1u1 + m2u2= m1v1+m2f2fm1 is the mass of bullet that is mm2 is the mass of wooden block that is M Initial Speed of bullet is v and that of wooden block is 0 m×v+M×0=m×y+M×ymv+0=my+Mymv=y (m+M)mv/m+M= yAnswer